When the mass of an electron becomes equal to thrice it's mass,its speed is

Grade:12

1 Answers

Arun
5 years ago
For a simple case, you could use the Lorenz equation 1/√(1 - v²/c²) which gives you a factor by which to multiply mass at relativistic velocities, where v = velocity and c = speed of light. 

For the mass to be 3 times rest mass, the factor must be 3. 
Therefore 
1/√(1 - v²/c²) = 3 
so 
v²/c² = 2/3 
v² = 2c²/3 

plug in constant c = 3 x 10^8 ms^-1 
c² = 9 x 10^16 
2c²/3 = 6 x 10^16 
Finally 
v² = 6 x 10^16 
so v = 245000000 m/s 
= just over 80% speed of light